OpenStudy (anonymous):

John is going to rent a tent to go camping. The cost to rent a camping tent from Woodland Outfitters is a flat rate of $12.00 plus $2.00 for each night the tent is rented. Part a. Let x represent the number of nights John keeps the tent and write a cost function, c(x), for the cost to rent a tent from Woodland Outfitters. Part b. How much will John owe if he uses the tent for three nights?

OpenStudy (blurbendy):

A) c(x)=2x+12 B) c(3)=2*3+12 =6+12 = $18